If rectangular room measures 13 feet by 19 feet, would the perimeter of the room be less than 30 yards
Answer by sofiyac(983)

P=2*13+2*19=26+38=64 feet, convert feet to yards....divide feet by 3
64/3=21.3333 yards
yes it will be less than 30 yards